OpenStudy (anonymous):

Do anybody know this? Show an equation and a solution for the problem. Arnold needs a 25% solution of nitric acid. He has 20 ml of a 30% solution. How many ml of a 15% solution should he add to obtain the required 25% solution?

OpenStudy (anonymous):

Let the amount of 15% solution required be = x ml ATQ xml of 15% + 20 ml of 30% = (x + 20) ml of 25% => x * 0.15 + 20*0.30 = (x + 20)*0.25 => 0.15x + 6 = 0.25x + 5 => 6 - 5 = 0.25x - 0.15x => 1 = 0.1x => 1/0.1 =x => x = 10 so 10 ml of 15% solution should he add to obtain the required 25% solution